# CoFiI2P: Coarse-to-Fine Correspondences-Based Image-to-Point Cloud Registration

### v3

@@ Abstract
Image-to-point cloud (I2P) registration is a fundamental task for robots and autonomous vehicles to achieve cross-modality data fusion and localization. Existing I2P registration methods estimate correspondences at the point/pixel level, often overlooking global alignment. However, I2P matching can easily converge to a local optimum when performed without high-level guidance from global constraints. To address this issue, this paper introduces CoFiI2P, a novel I2P registration network that extracts correspondences in a coarse-to-fine manner to achieve the globally optimal solution. First, the image and point cloud data are processed through a Siamese encoder-decoder network for hierarchical feature extraction. Second, a coarse-to-fine matching module is designed to leverage these features and establish robust feature correspondences. Specifically, In the coarse matching phase, a novel I2P transformer module is employed to capture both homogeneous and heterogeneous global information from the image and point cloud data. This enables the estimation of coarse super-point/super-pixel matching pairs with discriminative descriptors. In the fine matching module, point/pixel pairs are established with the guidance of super-point/super-pixel correspondences. Finally, based on matching pairs, the transform matrix is estimated with the EPnP-RANSAC algorithm. Extensive experiments conducted on the KITTI dataset demonstrate that CoFiI2P achieves impressive results, with a relative rotation error (RRE) of 1.14 degrees and a relative translation error (RTE) of 0.29 meters. These results represent a significant improvement of 84% in RRE and 89% in RTE compared to the current state-of-the-art (SOTA) method.   The project page is available at https://whu-usi3dv.github.io/CoFiI2P.
@@
Figure 1: Comparison of existing one-stage I2P registration scheme and proposed coarse-to-fine I2P registration scheme. (a) shows the one-stage registration pipeline, where matching pairs are directly established at the point/pixel level, leading to a significant number of mismatches. (b) shows our coarse-to-fine matching pipeline. Under the guidance of super point-to-pixel pairs, point-to-pixel pairs are generated from the existing super pairs, which effectively eliminates most mismatches.
@@ I Introduction
Estimating six degrees of freedom (6-DoF) pose of a monocular image concerning a pre-build point cloud map is a fundamental requirement for robots and autonomous vehicles  [1, 2]. Due to the limited onboard resources, robots are often equipped with only a monocular camera while facing challenges related to scale ambiguity in absolute localization and depth sensing. The establishment of a precise pose transformation between the image coordinate system and the pre-built point cloud coordinate system is important, as it not only accurately localizes the robot but also effectively reduces scale uncertainty inherent in monocular data [3].
However, cross-modality registration has its inherent challenges. Some existing methods use hand-crafted detectors and descriptors for I2P registration [4]. These approaches rely on structured features like edge features [5], which are limited by specific environmental conditions. With the rapid development of deep learning (DL), learning-based I2P registration approaches [6, 7, 8] have been proposed to extract representative keypoints and descriptors. However, the manually designed detectors for keypoints extraction lead to poor correspondence accuracy. To alleviate the difficulty of correspondence construction and improve the registration success rate, DeepI2P [7] converts the registration problem to a classification problem. A novel binary classification network is designed to distinguish whether the projected points are within or beyond the camera frustum. The classification results are passed into an inverse camera projection solver to estimate the transformation between the camera and laser scanners. As a large number of points on the boundary are misclassified, the accuracy of the camera pose is still limited. CorrI2P [8] proposed an overlap region detector for both image and point cloud, then pixels and points in the overlap region are matched to obtain I2P correspondences. Feature fusion module is exploited to fuse the point cloud and image information. Although CorrI2P [8] has significant improvement over DeepI2P [7], matching merely on one stage, namely, the pixel-point level without global alignment guidance, can lead to local minimal and instability.
Inspired by recent coarse-to-fine matching schedules and transformers in image-to-image (I2I) registration [9, 10] and point cloud-to-point cloud (P2P) registration approaches [11], this paper proposes the Coarse-to-Fine Image-to-Point cloud (CoFiI2P) network to tackle the challenges of existing one-stage I2P registration. I2P transformer with self-attention modules and cross-attention modules is embedded into the network for global alignment. CoFiI2P mainly contains four modules: feature extraction (FE), coarse matching (CM), fine matching (FM), and pose estimation (PE). FE module embeds input image and point cloud into shared high-dimensional spaces, and then CM and FM modules establish super-point/super-pixel correspondences and point/pixel correspondences. In the PE module, the relative transformation between the camera and laser scanners is estimated with EPnP-RANSAC algorithm [12, 13]. Overall, the main contributions of this paper are:
@@
    A novel coarse-to-fine I2P registration network is proposed to align image and point cloud in a progressive way. The coarse matching step provides robust but rough super-point/super-pixel correspondences for the fine matching step, which filters out most mismatched pairs and reduces the computation burden. The fine matching step achieves accurate and reliable point/pixel correspondences within the global guidance.
@@ II-A1 I2I Registration
before the age of deep-learning, hand-crafted detectors and descriptors (i.e., SIFT [14] and ORB [15]) are widely used to extract correspondences. Compared to traditional methods, learning-based methods improve the robustness and accuracy of image matching with large viewpoint differences and illumination changes. SuperPoint [16] proposed a self-learning training method through homography adaptation. SuperGlue [17] proposed an attention-based graph neural network (GNN) for feature matching. Patch2Pix [18] is the first work to obtain patch matches and regress pixel-wise matches in a coarse-to-fine manner. LoFTR [9] operates detector-free matching to get dense correspondences and overcomes small receive field of CNN with transformer. However, I2I registration methods can not be directly transferred to the I2P registration, which extracts heterogeneous features from cross-modality data.
@@ II-A2 P2P Registration
point cloud registration aims to estimate the optimal rigid transformation of two point clouds. Correspondence-based methods [19, 20] estimate the correspondence first and recover the transformation with robust estimation methods[13]. RoReg [21] embeds orientation information of point cloud to estimate local orientation and refine coarse rotation through residual regression to achieve fine registration. In the transformer era, point-based transformers [22, 11] have emerged and shown great performance. CoFiNet [22] followed LoFTR’s [9] design and proposed coarse-to-fine correspondences for registration, which computed coarse matches with descriptors strengthened by the transformer and refined coarse matches through density adaptive matching module. GeoTransformer [11] proposed a geometric Transformer to make full use of the 3D properties of the point cloud.
@@ II-B Cross-modality Registration
To address the cross-modality registration problem, a variety of I2P registration methods have been proposed, which could be roughly divided into two categories: I2P fine registration (initial transformation dependent) and I2P coarse registration (initial transformation free). The I2P fine registration methods [23, 4, 24] rely on initial transform parameters, and are widely applied in sensors calibration. Although this paper focuses on the second class, namely the coarse I2P registration without any initial transformation knowledge, we give an exhaustive review of both categories registration methods here.
@@ II-B1 Fine registration methods
fine registration methods have been thoroughly researched for several decades. Early-stage studies [25, 26] utilize various artificial targets as calibration constraints. In recent years, some approaches have argued that structured features shared in images and point cloud could be used for target-less I2P registration, e.g., edge feature [23] and line feature [4].  [23] employs edge information and optimizes transformation according to the response value.  [4] extracts lines from image and point clouds for registration. Recently, research on 2D and 3D semantic segmentation motivates semantic feature-based I2P registration. Se-calib [24] further studied semantic edges in I2P registration to employ more common semantic information instead of a certain type of object. Overall, fine registration methods achieve high accuracy but depend on the quality of the initial transformation value.
@@ II-B2 Coarse registration methods
2D3D-Matchnet [6] is the pioneering method to regress the relative transform parameters with CNN. It extracts SIFT [14] and ISS [27] keypoints from image and point cloud and learns descriptors with a Siamese network. However, the hand-crafted detectors from different modalities match poorly. DeepI2P [7] proposed a feature fusion module to merge image and point cloud information and classify points in/beyond the camera frustum. CorrI2P [8] predicts pixels and points in overlapping areas and matches with dense per-pixel/per-point features directly to get I2P correspondences. Although overlapping region detectors significantly reduce the number of false candidates, I2P registration only using the low-level feature without global guidance leads to serious mismatches. Inspired by the coarse-to-fine strategies in CoFiNet [22], we propose the coarse-to-fine I2P registration network CoFiI2P, which injects the information of high-level correspondences into low-level matching for rejecting mismatches.
@@ III Methodology
Figure 2: Workflow of CoFiI2P. The proposed method consists of feature extraction, coarse matching, fine matching and pose estimation modules. Image and point cloud are sent to the feature extraction module to obtain hierarchical deep features, respectively. The coarse-level features are strengthened by I2P transformer module and then matched with the cosine similarity rule. Fine features are gathered from the last layer of the decoder. In each super-point/super-pixel pair, the node point is set as the candidate and the corresponding pixel is selected from the super-pixel area, a $s\times s$ window. The generated dense matching pairs are utilized to regress the pose with the EPnP-RANSAC [12, 13] algorithm.
@@
For convenient description, a pair of partially overlapped image and point cloud are defined as $\mathrm{I}\in\mathbb{R}^{W\times H\times 3}$ and $\mathrm{P}\in\mathbb{R}^{N\times 3}$, where $W$ and $H$ are the width and height, and $N$ is the number of points. The purpose of I2P registration is to estimate the relative transformation between the image $\mathrm{I}$ and point cloud $\mathrm{P}$, including rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and translation vector $\mathbf{t}\in\mathbb{R}^{3}$.
@@
Our method adopts the coarse-to-fine manner to find the correct correspondences set $\Omega(\mathrm{I},\mathrm{P})$. The CoFiI2P mainly consists of four modules: feature extraction (FE), coarse matching (CM), fine matching (FM) and pose estimation (PE). FE is an encoder-decoder structure network, that encodes raw inputs from different modalities into high-dimension feature spaces and finds in-frustum super-points. CM and FM are cascaded two-stage matching modules. CM constructs coarse matching at super-pixel/super-point level, and them FM constructs fine matching at pixel/point level sequentially with the guidance of super-pixel/super-point correspondences. Lastly, the PE module exploits point-pixel matching pairs to regress the relative pose with EPnP-RANSAC [12, 13] algorithm. The workflow of the proposed method is shown in Fig. 2.
III-A Multi-scale Feature extraction
@@
We utilize ResNet-34 [28] and KPConv-FPN [29] as the backbones for image and point cloud to extract multi-level features. The encoder progressively embeds raw inputs into high-dimensional features, and the decoder propagates high-level information to low-level details with skip-connection for dense per-pixel/point feature generation. Then, we extract features from multiple resolutions for coarse-to-fine matching. Specifically, super-points set $\tilde{P}$ and super-pixels set $\tilde{I}$ at the coarsest resolution are chosen as candidates for coarse matching, and local points group $G_{\tilde{p}}$ and pixels patches $G_{\tilde{i}}$ are generated from coarse matching pairs for fine matching. The local points group $G_{\tilde{p}_{x}}$ are constructed with the point-to-node strategy in the geometric space:
$G_{\tilde{p}_{x}}=\{p\in P\lvert\|p-\tilde{p}_{x}\|<r_{g}\},$ (1)
where $r_{g}$ is the chosen radius. As the pixels array on the image with a rigid order, the local pixels patch $G_{\tilde{i}}$ construct simply with the pyramid matching strategy.
@@ III-B1 I2P Transformer
$\mathbf{q}=\mathbf{W_{q}F},\mathbf{k}=\mathbf{W_{k}F},\mathbf{q}=\mathbf{W_{v}F},$ (2)
@@
Extracted global-aware features $\mathbf{F}^{sa}$ are fed into the feed-forward network (FFN) to fuse the spatial relation information in channel dimension. Given a feature map $\mathbf{F}$, the relative positions are encoded with multi-layer perception (MLP) [33].
@@
where $\mathbf{q}_{\tilde{P}},\mathbf{k}_{\tilde{P}},\mathbf{v}_{\tilde{P}}$ are query, key and value vectors of point cloud feature $\mathbf{F}_{\tilde{P}}$, and $\mathbf{q}_{\tilde{I}},\mathbf{k}_{\tilde{I}},\mathbf{v}_{\tilde{I}}$ are query, key and value vectors of image feature $\mathbf{F}_{\tilde{I}}$.
Remark. While the self-attention module encodes the spatial and geometric features for each super-pixel and super-point, the cross-attention module injects the geometric structure information and texture information across image and point cloud respectively. Outputs of the I2P transformer carry powerful cross-modality information for matching.
III-B2 Super-point/Super-pixel Matching
for the monocular camera, the field of view (FoV) is obviously smaller than the laser scans of 3D Lidar (e.g., Velodyne-H64), which usually sweeps 360 degrees in the horizontal direction. Therefore, only a small number of super-points are in the camera frustum. To filter out beyond-frustum super-points, we add a simple binary classification head to predict super-points in or beyond the frustum of the camera. After the beyond-frustum super-points are removed, the coarse level correspondences are estimated between in-frustum super-points set $\tilde{P}$ and super-pixels set $\tilde{I}$ by sorting the nearest super-pixel $\tilde{i}$ in feature space:
$\begin{split}\tilde{\Omega}_{\mathcal{M}}=\{(\tilde{p}_{x}\in\tilde{P},\tilde{i}_{y}\in\tilde{I})\lvert y=\mathop{\mathrm{argmin}}\limits_{\lvert\tilde{I}\rvert}\|\mathbf{F}_{\tilde{P}}(\tilde{p}_{x})-\mathbf{F}_{\tilde{I}}(\tilde{i}_{y})\|\},\end{split}$ (5)

### v4

# CoFiI2P: Coarse-to-Fine Correspondences-Based Image to Point Cloud Registration

@@ Abstract
Image-to-point cloud (I2P) registration is a fundamental task for robots and autonomous vehicles to achieve cross-modality data fusion and localization. Current I2P registration methods primarily focus on estimating correspondences at the point or pixel level, often neglecting global alignment. As a result, I2P matching can easily converge to a local optimum if it lacks high-level guidance from global constraints. To improve the success rate and general robustness, this paper introduces CoFiI2P, a novel I2P registration network that extracts correspondences in a coarse-to-fine manner. First, the image and point cloud data are processed through a two-stream encoder-decoder network for hierarchical feature extraction. Second, a coarse-to-fine matching module is designed to leverage these features and establish robust feature correspondences. Specifically, In the coarse matching phase, a novel I2P transformer module is employed to capture both homogeneous and heterogeneous global information from the image and point cloud data. This enables the estimation of coarse super-point/super-pixel matching pairs with discriminative descriptors. In the fine matching module, point/pixel pairs are established with the guidance of super-point/super-pixel correspondences. Finally, based on matching pairs, the transform matrix is estimated with the EPnP-RANSAC algorithm. Experiments conducted on the KITTI Odometry dataset demonstrate that CoFiI2P achieves impressive results, with a relative rotation error (RRE) of 1.14 degrees and a relative translation error (RTE) of 0.29 meters, while maintaining real-time speed. These results represent a significant improvement of 84% in RRE and 89% in RTE compared to the current state-of-the-art (SOTA) method. Additional experiments on the Nuscenes datasets confirm our method’s generalizability. The project page is available at https://whu-usi3dv.github.io/CoFiI2P.
@@
Figure 1: Comparison of existing one-stage I2P registration and proposed coarse-to-fine I2P registration. (a) The existing one-stage registration pipeline. The matching pairs are directly established at the point/pixel level, leading to a significant number of mismatches. (b) Our coarse-to-fine matching pipeline. Under the guidance of super point-to-pixel pairs, point-to-pixel pairs are generated from the existing super pairs, which effectively eliminates most mismatches.
@@ I Introduction
Estimating the six degrees of freedom (6-DoF) pose of a monocular image relative to a pre-built point cloud map is a fundamental requirement for robots and autonomous vehicles. [1, 2]. Due to the limited onboard resources, robots are often equipped with only a monocular camera while facing challenges related to scale ambiguity in absolute localization and depth sensing. Establishing an accurate pose transformation between the image coordinate system and the pre-built point cloud coordinate system is crucial. This transformation not only precisely localizes the robot but also effectively reduces the scale uncertainty inherent in monocular data [3].
However, cross-modality registration has its inherent challenges. Some existing methods use hand-crafted detectors and descriptors for I2P registration [4, 5]. These approaches rely on structured features like edge features [5], which are limited by specific environmental conditions. With the rapid development of deep learning (DL), learning-based I2P registration approaches [6, 7, 8] have been proposed to extract representative keypoints and descriptors. 2D3D-MatchNet [6] proposed a novel Siamese network to learn the cross-modality descriptor but the manually designed detectors matched poorly. To alleviate the difficulty of correspondence construction and improve the registration success rate, DeepI2P [7] converts the registration problem to a classification problem. A novel binary classification network is designed to distinguish whether the projected points are within or beyond the camera frustum. The classification results are passed into an inverse camera projection solver to estimate the transformation between the camera and laser scanners. As a large number of points on the boundary are misclassified, the accuracy of the camera pose is still limited. CorrI2P [8] proposed an overlap region detector for both image and point cloud, then pixels and points in the overlap region are matched to obtain I2P correspondences. Feature fusion module is exploited to fuse the point cloud and image information. Although CorrI2P [8] has significant improvement over DeepI2P [7], matching merely on one stage, namely, the pixel-point level without global alignment guidance, can lead to local minima and instability.
Inspired by recent coarse-to-fine matching schedules and transformers in image-to-image (I2I) registration [9, 10] and point cloud-to-point cloud (P2P) registration approaches [11], this paper proposes the Coarse-to-Fine Image-to-Point cloud (CoFiI2P) network to for I2P registration. I2P transformer with self- and cross-attention modules is embedded into the network for global alignment.
Overall, the main contributions of this work are as follows.
@@
    A novel coarse-to-fine I2P registration network is proposed to align image and point cloud in a progressive way. The coarse matching step provides rough but robust super-point/super-pixel correspondences for the following fine matching step, which filters out most mismatched pairs and reduces the computation burden. The fine matching step achieves accurate and reliable point/pixel correspondences with the global guidance.
@@ II-A1 I2I Registration
Before the age of deep-learning, hand-crafted detectors and descriptors (i.e., SIFT [12] and ORB [13]) are widely used to extract correspondences for matching. Compared to traditional methods, learning-based methods improve the robustness and accuracy of image matching with large viewpoint differences and illumination changes. SuperPoint [14] proposed a self-learning training method through homography adaptation. SuperGlue [15] proposed an attention-based graph neural network (GNN) for feature matching. Patch2Pix [16] is the first work to obtain patch matches and regress pixel-wise matches in a coarse-to-fine manner.
@@ II-A2 P2P Registration
Point cloud registration aims to estimate the optimal rigid transformation of two point clouds. Correspondence-based methods [17, 18] estimate the correspondence first and recover the transformation with robust estimation methods [19]. RoReg [20] embeds orientation information of point cloud to estimate local orientation and refine coarse rotation through residual regression to achieve fine registration. In the transformer era, point-based transformers [21, 11] have emerged and shown great performance. CoFiNet [21] followed LoFTR’s [9] design and proposed coarse-to-fine correspondences for registration, which computed coarse matches with descriptors strengthened by the transformer and refined coarse matches through density adaptive matching module. GeoTransformer [11] uses Transformer to fully exploit the 3D properties of point clouds. However, I2I or P2P registration methods can not be directly used for the I2P registration, which extracts heterogeneous features from cross-modality data.
@@ II-B Cross-modality Registration
To address the cross-modality registration problem, a variety of I2P registration methods have been proposed, which could be roughly divided into two categories: I2P fine registration (initial transformation dependent) and I2P coarse registration (initial transformation free). The I2P fine registration methods [22, 4, 23] rely on initial transform parameters and are widely applied in sensor calibration. Although this paper focuses on the second category, namely coarse I2P registration without any initial transformation knowledge, we provide a comprehensive review of registration methods in both categories.
@@ II-B1 Fine Registration Methods
Fine registration methods have been thoroughly studied for several decades. Early-stage works [24, 25] utilize various artificial targets as calibration constraints. In recent years, some approaches have argued that structured features shared in images and point cloud could be used for target-less I2P registration, e.g., edge feature [22] and line feature [4][22] employs edge information and optimizes transformation according to the response value. [4] extracts lines from image and point clouds for registration. Recently, research on 2D and 3D semantic segmentation motivates semantic feature-based I2P registration. Se-calib [23] further studied semantic edges in I2P registration to employ more common semantic information instead of a certain type of object. Overall, fine registration methods achieve high accuracy but depend on the quality of the initial values.
@@ II-B2 Coarse Registration Methods
2D3D-Matchnet [6] is the pioneering method to regress the relative transform parameters with CNN. It extracts SIFT [12] and ISS [26] keypoints from image and point cloud and learns descriptors with a Siamese network. Experimental results show that hand-crafted detectors from different modalities match poorly. HAS-Net [27] proposed a novel network for learning cross-domain descriptors from 2D image patches and 3D point cloud volumes. However, both 2D3D-Matchnet and HAS-Net split the image and point cloud into patches and volumes, and then match with the descriptors of patches and volumes, resulting in the loss of long-range context and high-level information. DeepI2P [7] proposed a feature fusion module to merge image and point cloud information and classify points in/beyond the camera frustum. CorrI2P [8] predicts pixels and points in overlapping areas and matches with dense per-pixel/per-point features directly to get I2P correspondences. Although overlapping region detectors significantly reduce the number of false candidates, I2P registration only using the low-level feature without global guidance leads to serious mismatches. Inspired by the coarse-to-fine strategy in CoFiNet [21], we propose the CoFiI2P network, a coarse-to-fine I2P registration approach that integrates high-level correspondence information into low-level matching to effectively reject mismatches.
@@ III Methodology
Figure 2: Workflow of CoFiI2P. The proposed method consists of feature extraction, coarse matching, fine matching and pose estimation modules. Image and point cloud are sent to the feature extraction module to obtain coarse-level features and fine-level features (rendered in red for image and green for point cloud, respectively). The coarse-level features are strengthened by I2P transformer module and then matched with the cosine similarity. Fine features are gathered from the last layer of the decoder. In each super-point/super-pixel pair, the node point is set as the candidate and the corresponding pixel is selected from the super-pixel area, a $s\times s$ window. The generated fine-level matching pairs are utilized to estimate the pose with the EPnP-RANSAC [28, 19] algorithm.
@@
For convenient description, a pair of partially overlapped image and point cloud are defined as $\mathrm{I}\in\mathbb{R}^{W\times H\times 3}$ and $\mathrm{P}\in\mathbb{R}^{N\times 3}$, where $W$ and $H$ are the width and height, and $N$ is the number of points. The purpose of I2P registration is to estimate the relative transformation between the image $\mathrm{I}$ and point cloud $\mathrm{P}$, defined by a rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and a translation vector $\mathbf{t}\in\mathbb{R}^{3}$.
@@
Our method adopts the coarse-to-fine manner to find the correct correspondences set $\Omega(\mathrm{I},\mathrm{P})$ and calculate the relative pose using EPnP-RANSAC [28]. The CoFiI2P mainly consists of four modules: feature extraction (FE), coarse matching (CM), fine matching (FM), and pose estimation (PE). FE is an encoder-decoder structure network, that encodes raw inputs from different modalities into shared feature spaces and finds in-frustum super-points. CM and FM are cascaded two-stage matching modules. CM constructs coarse matching at super-pixel/super-point level, and them FM constructs fine matching at pixel/point level sequentially with the guidance of super-pixel/super-point correspondences. Lastly, the PE module exploits point-pixel matching pairs to regress the relative pose with the EPnP-RANSAC algorithm. The workflow of the proposed method is shown in Fig. 2.
III-A Feature Extraction
@@
We utilize ResNet-34 [29] and KPConv-FPN [30] as the backbones for image and point cloud to extract multi-level features. The encoder progressively embeds raw inputs into high-dimensional features, and the decoder propagates high-level features to low-level features with skip-connection for per-pixel/point feature generation. Specifically, points and pixels at the coarsest resolution ($\frac{1}{8}$ of original resolution) ${\tilde{P}}$ and ${\tilde{I}}$ are treated as superpoints and superpixels for coarse matching, and ${P}$ and ${I}$ at the finest resolution ($\frac{1}{2}$ of original resolution) are used for fine matching. We use $\boldsymbol{F}_{\tilde{P}}$ and $\boldsymbol{I}_{\tilde{P}}$ to denote coarse-level features, and $\boldsymbol{F}_{P}$ and $\boldsymbol{I}_{P}$ for fine-level features. For each superpoint, we construct the local points group $G_{\tilde{p}}$ with point-to-node strategy in the geometric space:
$G_{\tilde{p}}=\{p\in P\lvert\|p-\tilde{p}\|<r_{g}\},$ (1)
where $r_{g}$ is the chosen radius. For the superpixels, we first locate their positions on the fine-level feature map, and then crop sets of local pixel patches with window size $w\times w$.
@@ III-B1 I2P Transformer
$\mathbf{q}=\mathbf{W_{q}F},\mathbf{k}=\mathbf{W_{k}F},\mathbf{v}=\mathbf{W_{v}F},$ (2)
@@
Softmax operates row-wise on the attention matrix $\mathbf{q}\mathbf{k}^{\top}$ to obtain the weights on the values. Extracted global-aware features $\mathbf{F}^{sa}$ are fed into the feed-forward network (FFN) to fuse the spatial relation information in channel dimension. Given a feature map $\mathbf{F}$, the relative positions are encoded with multi-layer perception (MLP) [34].
@@
where $\mathbf{q}_{\tilde{P}},\mathbf{k}_{\tilde{P}},\mathbf{v}_{\tilde{P}}$ are query, key and value vectors of point cloud feature $\mathbf{F}_{\tilde{P}}$, and $\mathbf{q}_{\tilde{I}},\mathbf{k}_{\tilde{I}},\mathbf{v}_{\tilde{I}}$ are query, key and value vectors of image feature $\mathbf{F}_{\tilde{I}}$. The softmax operation is the same as in the self-attention module.
Remark 1. While the self-attention module encodes the spatial and geometric features for each super-pixel and super-point, the cross-attention module injects the geometric structure information and texture information across image and point cloud respectively. Outputs of the I2P transformer carry powerful cross-modality information for matching.
III-B2 Super-point/-pixel Matching
The field of view (FoV) of monocular cameras is obviously smaller than the laser scans of 3D Lidar (e.g., Velodyne-H64), which sweeps 360 degrees in the horizontal direction. Therefore, only a small number of super-points are in the camera frustum. To filter out beyond-frustum super-points, we add a simple binary classification head to predict super-points in or beyond the frustum of the camera. The coarse level correspondences are estimated between in-frustum super-points set $\tilde{P}$ and super-pixels set $\tilde{I}$ by finding the nearest super-pixel $\tilde{i}$ in feature space:
$\tilde{\Omega}_{\mathcal{M}}=\{(\tilde{p}_{x}\in\tilde{P},\tilde{i}_{y}\in\tilde{I})\lvert y=\mathop{\mathrm{argmin}}\limits_{\lvert\tilde{I}\rvert}\|\mathbf{F}_{\tilde{P}}(\tilde{p}_{x})-\mathbf{F}_{\tilde{I}}(\tilde{i}_{y})\|\},$ (5)